OpenStudy (anonymous):

I need help with my math word problem. Question: SAra pays 3700 for tution and fees for each of the two semesters, plus additional $440 for textbooks each semester. what is the prorated monthly cost for tutition and fees and extbooks?

OpenStudy (ghazi):

3700*2= $ 7400 for 12 months...therefore for one month 7400/12= $616.67 and on books 440*2=$ 880 for 12 months therefore for one month =$ 73.33....therefore total monthly expense= 73.33+616.67= $690